OpenStudy (anonymous):

a sweater originally priced at $32 is on sale for $28.what is the percent of decrease rounded to the nearest whole number?

OpenStudy (jmark):

Given original priced =$32 on sale=$28 Decrease price =32-28=$4 decrease in %=(original price-sale price)*100/(original price) percent of decrease =(32-28)*100/32 percent of decrease =400/32 percent of decrease=12.5% ICSE 8th Standard Question Papers - http://education-help-online.blogspot.in/2014/07/icse-8th-standard-question-papers.html